A moving object A of mass 8.7 kg and speed 2.2 m/s collides with a stationary object B of mass 9.4 kg. The collision is totally inelastic. Calculate the speed of the combined object after collision.
Answer #1

The collision is totally inelastic so the two bodies stick together.

Applying conservation of momentum we have

Initial momentum = final momentum

So, 8.7×2.2 =( 8.7+9.4) v

Thus , v = (8.7×2.2)/ 18.1 = 1.06 m/s.